OpenStudy (anonymous):

What is the value of the expression? –13.7 + (–5 + 2.7) A. –21.4 B. –16 C. –11.4 D. –6

OpenStudy (adamaero):

2-5 = ?

OpenStudy (anonymous):

-3

OpenStudy (adamaero):

-5+2.7 = ?

OpenStudy (anonymous):

-2.3

OpenStudy (anonymous):

-16

OpenStudy (anonymous):

–13.7 + (–5 + 2.7) -13.7+(-5+2.7) =-13.7+(-2.3) =-16

OpenStudy (anonymous):

ok tyvm

OpenStudy (anonymous):

you are welcome @hurica1n :)